OpenStudy (anonymous):

sqrt(2*m^2*n^3) * sqrt(20*n)

OpenStudy (anonymous):

just multiply them

OpenStudy (jhannybean):

\[\large \sqrt{2m^2n^3}\cdot \sqrt{20n}\]\[\large \sqrt{40m^2n^4}\]\[\large \sqrt{5\cdot \color{green}{2\cdot 2\cdot 2}\color{purple}{\cdot m\cdot m}\cdot\color{blue}{ n \cdot n \cdot n\cdot n}}\] for every par of 2 of these numbers,you can pull one out, everything else gets left inside. \[\large 2\cdot m\cdot n^2\sqrt{10}\]